OpenStudy (anonymous):

How can we solve \[ \lim_{h \rightarrow 0}{ \frac{\|x+h\| - \|x\|}{h} } \]

OpenStudy (anonymous):

h -> 0 but 0^- or 0^+??

OpenStudy (anonymous):

just 0

OpenStudy (anonymous):

The answer is \[ \|x\|/x \] because that is \[\frac{d}{dx} \|x\| \] But not sure how I can get there with limits

OpenStudy (anonymous):

0

OpenStudy (anonymous):

It's not 0 the derivative on the real axis is defined in any point except for 0

OpenStudy (shubhamsrg):

lets use definition of |x| |x| = sqrt(x^2) making that substitution, we have sqrt((x+h)^2) - sqrt(x^2) /h on rationalizing , [ (x+h)^2 - x^2 ]/h (sqrt((x+h)^2) + sqrt(x^2)) (x+h -x)(x+h+x) /h (sqrt((x+h)^2) + sqrt(x^2)) => (2x+h)/h (sqrt((x+h)^2) + sqrt(x^2)) now put h=0 => 2x / 2sqrt(x^2) => x/|x| hmm,,this was a good question..

OpenStudy (shubhamsrg):

Correction : in the 3rd last step , h won't be there in the denominator, it gets cancelled from the numerator.

OpenStudy (anonymous):

how you get 2x+h all of a sudden?

OpenStudy (anonymous):

@shubhamsrg

Parth (parthkohli):

That's just the definition of the derivative for \(||x||\).

OpenStudy (anonymous):

It is, but the question was how to solve it with limits

OpenStudy (anonymous):

No it's |x|/x as I stated aboved and it's defined for x =/= 0 in R

zepdrix (zepdrix):

@xcrypt Shrub showed you the steps. Did you get confused by any of that?

OpenStudy (anonymous):

I don't know why he got 2x+h in the numerator all of a sudden

zepdrix (zepdrix):

\[f(x)=|x|\]We're going to write it like this instead, \(f(x)=\sqrt{x^2}\) So for \(f(x+h)\) we have, \(\sqrt{(x+h)^2}\) So we'll plug these pieces into our derivative definition.\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h} \qquad =\qquad \lim_{h \rightarrow 0}\frac{\sqrt{(x+h)^2}-\sqrt{x^2}}{h}\]This is the confusing part I guess :) You want to rationalize the top. So multiply the top and bottom by the CONJUGATE of the top. Understand how to do that?

zepdrix (zepdrix):

\[\lim_{h \rightarrow 0}\frac{\sqrt{(x+h)^2}-\sqrt{x^2}}{h}\left(\frac{\sqrt{(x+h)^2}+\sqrt{x^2}}{\sqrt{(x+h)^2}+\sqrt{x^2}}\right)\]

zepdrix (zepdrix):

The top will end up simplifying to 2xh+h^2, which will then further simplify to 2x+h. If you're having trouble getting through the simplification let me know.

OpenStudy (shubhamsrg):

Well I used a^2 - b^2 = (a+b)(a-b) that will yield you (2x+h)(h) also |x|/x = x/|x| both yield -1 when x<0 and +1 when x>0